OpenStudy (anonymous):

given that one zero is -4 find all zeros of p(x)=x^3+x^2-22x-40

OpenStudy (anonymous):

i got 5,-4, and -2

OpenStudy (solomonzelman):

yes, correct

OpenStudy (anonymous):

thank you :3

OpenStudy (solomonzelman):

All I did is checked via wolfram, not much, but you are welcome in any case :)

jimthompson5910 (jim_thompson5910):

Here's a way to check If you claim that `-4 is a zero` then plugging in x = -4 should make p(x) = 0 Let's check p(x)=x^3+x^2-22x-40 p(-4)=(-4)^3+(-4)^2-22(-4)-40 ... replace every x with -4 p(-4)=-64+16+88-40 p(-4)=0 since the result is 0, the claim `-4 is a zero` is true You would test the other roots the same way